if a body is droped from 100m it takes t seconds to reach the ground.then find its height at time=t\2 seconds......

Grade:10

3 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Aditya

take downward direction as positive

Y=uyt + at2/2

put Y=100m, a=g=10 m/s2, uy=0

100=5t2

t=2√5 sec

now t/2=√5 sec

distance covered after t/2 sec

again apply the formula

Y=uyt + at2/2

and put t=√5 sec

Y=10*√5*√5/2

Y=25m

height from ground= 100-25=75m

 

All the best.

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Amit Tapas
13 years ago
See, by 2 kinematical equation s=ut+1/2 at(squared) given s=100 t=t u=0 therefore 100=1/2*a*t(squared) so a=200/t(squared) to find height when t=t/2 s=ut+1/2at(squared) s=1/2*200/t(squared)*t(squared)/4 s=25m the body reaches the height of 25m at t/2 seconds
vikas askiitian expert
13 years ago

body takes t sec to reach the groung..

initial velocity is zero..

s=ut +at2 /2

-100=-gt2 /2

 t=2sqrt5 sec

now displacement of particle=s=gT2 /2

for T=t/2=sqrt5sec

 S=25m

so the particle is at (100-25)=75m above the ground at T=t/2...